# LiDAR, GNSS and IMU Sensor Alignment through Dynamic Time Warping to Construct 3D City Maps

### v2

@@ III-C High-Precision 3D LiDAR Map Construction
Figure 6 illustrates the proposed mapping pipeline. It comprises three main stages: (1) loop registration using the NDT algorithm and pose graph optimization with loop closures; (2) GNSS-based calibration to anchor local trajectories to geographic coordinates; and (3) global merging of loop segments via ICP alignment on overlapping regions. Each stage is detailed in the following subsections.
@@ III-C2 GNSS-Based Loop Calibration
Here, $\mathbf{W}_{i}$ is a diagonal weighting matrix that reflects the quality of the GNSS signal, including factors such as position dilution of precision (PDOP) and the status of the RTK fix. This step anchors local maps to real-world coordinates while de-emphasizing GNSS-denied segments with low weight, thereby preserving local loop integrity.
@@ III-C3 Global Merging Using ICP on Overlapping Regions
Final map integration is performed by aligning adjacent loops using ICP on overlapping point cloud regions. Fast Point Feature Histograms (FPFH) guide initial correspondence estimation, while Octree-based indexing identifies candidate overlaps. The alignment process minimizes the mean squared error between corresponding point pairs $(\mathbf{a}_{k},\mathbf{b}_{k})$ from each loop. This is formalized by the cost function $E(\mathbf{R},\mathbf{t})=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{R}\mathbf{a}_{k}+\mathbf{t}-\mathbf{b}_{k}\right\|^{2}$, where $\mathbf{R}$ and $\mathbf{t}$, where $\mathbf{R}$ and $\mathbf{t}$ denote the rotation and translation applied to the source point cloud. To improve robustness in dynamic environments (e.g., moving vehicles, pedestrians), a Trimmed ICP variant discards high-residual point pairs. Post-alignment, the map is refined using voxel grid downsampling (0.05 m resolution) and statistical outlier removal, enhancing compactness and structural clarity. The final map provides a reliable basis for downstream use.
@@ III-D2 Centreline-Based Alignment Error Metrics
Figure 8: Velocity profiles from IMU (top, blue) and LiDAR (bottom, red) before synchronization. Shaded regions indicate detected stationary phases. Misalignment across modalities underscores the need for temporal correction.
@@ III-D3 Intersection-Based Local Geometric Consistency Metrics
$\mathcal{E}_{\text{spatial}}=\frac{1}{K}\sum_{i=1}^{K}\left[\alpha\|p_{i}-m_{i}\|_{2}+\beta\cdot d_{\mathcal{T}}(p_{i},m_{i})\right]\vskip-5.69054pt$ (7)
@@
$\mathcal{E}_{\text{env}}=\frac{1}{K}\sum_{i=1}^{K}\max_{x\in\mathcal{B}(p_{i},r)}\min_{y\in\mathcal{B}(m_{i},r)}\|x-y\|_{2}\vskip-5.69054pt$ (8)
@@ IV Results
This section evaluates the performance of the proposed framework through a four-stage analysis. First, the experimental setup is outlined to support reproducibility. Second, sensor synchronization accuracy is examined, with emphasis on temporal alignment across LiDAR, GNSS, and IMU streams. Third, the results of 3D map reconstruction for Perth CBD are benchmarked against LiDAR-only baselines. Finally, skeleton-based centreline metrics are employed to assess spatial accuracy and geometric consistency. Each stage is described in detail in the subsequent subsections.
@@ IV-A Experimental Setup
The experiments were conducted on a high-performance workstation equipped with an Intel Xeon Gold 6330 CPU (28 cores), 512 GB of RAM, and an NVIDIA RTX A6000 GPU, running Ubuntu 20.04 LTS. To ensure reproducibility, all software components were containerized using Docker. The data processing pipeline utilized the Ouster SDK v3.1.1 for LiDAR decoding, Open3D v0.17 for point cloud filtering, and the Inertial Sense uINS SDK v2.15 for GNSS–IMU data handling. Pose graph optimization was performed using GTSAM 4.2, while NDT and ICP registration were implemented via the Point Cloud Library (PCL) v1.13.
@@ IV-B Effectiveness of Multimodal Sensor Data Synchronization
We evaluated the proposed synchronization framework by comparing IMU and LiDAR velocity profiles before and after alignment. Accurate temporal calibration is critical for sensor fusion, especially for pose estimation and 3D mapping. Despite hardware-level synchronization via PPS triggers, residual time offsets often persist due to jitter, latency, or clock drift, which results in misaligned motion events, most evident in dynamic urban environments. As shown in Figure 8, pre-alignment profiles display mismatched transitions and inconsistent stationary phases, degrading point cloud registration and pose accuracy. To address this, we applied a DTW-based velocity matching method, which automatically estimated a $1.45\,\mathrm{s}$ offset. Post-synchronization results (Figure 9) demonstrate close alignment of both static and dynamic phases, effectively synchronizing acceleration and deceleration events across sensors. Quantitatively, the IMU–LiDAR velocity cross-correlation exceeds $0.97$, and the RMSE drops below $0.15\,\mathrm{m/s}$, confirming high-fidelity temporal alignment. This robust synchronization enables reliable sensor fusion and supports downstream tasks, especially in stop-and-go traffic scenarios such as those observed in Perth CBD.
@@ IV-D1 Visual Comparison of Mapping Results
TABLE I: Comparison of road centreline alignment metrics across all test sequences for the LiDAR-only baseline and the proposed framework. Relative improvements indicate percentage reductions in alignment error.
@@ IV-D2 Loop-Based Centreline Comparison
To assess route-level consistency, we selected six representative loops from the twenty-one sequences in the Perth CBD (Seq 05, 11, 14, 16, 19, and 20). These loops span long straight segments, curved roads, and dense junctions. Figure 11 consolidates the visual evidence: for each loop, the top row shows the LiDAR-only baseline overlaid on satellite imagery, the middle row shows the reconstruction produced by LIGMA, and the bottom row presents the corresponding road centreline comparison against a 2D cartographic reference. The baseline typically exhibits drift-induced shear and irregular loop closures, with noticeable misalignment at intersections and along extended segments. In contrast, LIGMA yields coherent loop geometry and improved intersection alignment, with centrelines closely following the reference. While Fig. 11 illustrates typical behavior on six loops, the quantitative evaluation over all twenty-one sequences is summarized in Table I: our framework reduces centreline RMSE from $3.32$ m (LiDAR-only) to $1.24$ m on average, with consistent gains in the minimum, maximum, and standard deviation of the alignment error.
TABLE II: Intersection-level alignment metrics for 15 of 51 intersections (subset shown due to space constraints). CI: Centroidal Offset (m), MOD: Mean Orientation Deviation (°). Abs. = absolute reduction; Pct. = percentage reduction relative to baseline.
@@
TABLE III: Summary statistics of centreline and intersection alignment metrics averaged over all test sequences. CI: Centroidal Offset (m); MOD: Mean Orientation Deviation (°); RMSE: Root Mean Square Error of centreline deviation (m); SD: Standard Deviation (m). Relative improvements are computed with respect to the LiDAR-only baseline.
@@
IV-D3 Intersection-Based Alignment Evaluation
Table II summarizes results for 15 selected intersections, with percentage improvements shown in the right most columns. Table III summarizes average centerline and intersection alignment metrics across all test sequences. The proposed approach reduces centroid offset by 61.5%, orientation deviation by 55.1%, and LEE by 59.4% over the baseline. These improvements are most pronounced at challenging junctions and map boundaries. Accurate intersection alignment supports critical tasks such as trajectory planning, and robust localization. These results demonstrate the effectiveness of the proposed GNSS-constrained fusion framework.

### v3

@@ III-C High-Precision 3D LiDAR Map Construction
Figure 6 illustrates the proposed mapping pipeline. It comprises three main stages: (1) loop registration using the NDT algorithm and pose graph optimization with loop closures; (2) GNSS-based calibration to anchor local trajectories to geographic coordinates; and (3) global merging of loop segments via ICP alignment on overlapping regions. Since ICP can be unreliable under weak initialization, small overlap, dynamic objects and repeated structures, we invoke ICP only as a refinement step after GNSS-based anchoring.  Moreover, we apply strict overlap validation to ensure that there is sufficient overlap between the two point clouds. We also perform robust correspondence trimming using distance gating and geometric consistency checks to reduce sensitivity to moving objects and outliers. Finally, in case of a failed refinement, our method performs re-initialization using  global registration on feature correspondences and then retries the refinement, providing a practical recovery mechanism in difficult overlap conditions. Each stage is detailed in the following subsections.
@@ III-C2 GNSS-Based Loop Calibration
Here, $\mathbf{W}_{i}$ is a diagonal weighting matrix that reflects the quality of the GNSS signal, including factors such as position dilution of precision (PDOP) and the status of the RTK fix. This step anchors local maps to real-world coordinates while de-emphasizing GNSS-denied segments with low weight, thereby preserving local loop integrity. When a $3\times 3$ GNSS position covariance $\boldsymbol{\Sigma}_{i}$ is available, we set $\mathbf{W}_{i}=\bigl(\boldsymbol{\Sigma}_{i}+\sigma_{0}^{2}\mathbf{I}_{3}\bigr)^{-1}$ with $\sigma_{0}=0.05\,\text{m}$ as a stability floor. If covariance is unavailable, we use a mode-dependent diagonal model $\mathbf{W}_{i}=\operatorname{diag}(\sigma_{E,i}^{-2},\sigma_{N,i}^{-2},\sigma_{U,i}^{-2})$ with $(\sigma_{E,i},\sigma_{N,i},\sigma_{U,i})=(0.03,0.03,0.05)\,\text{m}$ for RTK FIX and $(0.50,0.50,1.00)\,\text{m}$ otherwise. Epochs with PDOP>6>6 or missing RTK status are discarded.
GNSS samples $\mathbf{g}_{i}$ are time-aligned to $\mathbf{p}_{i}$ by linearly interpolating the GNSS stream and pairing each pose with the nearest timestamp within $50$ ms; pairs outside this tolerance are discarded. In this anchoring step we align only the translation $\mathbf{p}_{i}$ and keep the loop’s internal rotations unchanged to preserve local loop shape.
For the least squares representation, let $\mathbf{p}_{i}\in\mathbb{R}^{3}$ be the translation of $\mathbf{T}_{\text{loop}}^{(i)}$ and $\mathbf{g}_{i}\in\mathbb{R}^{3}$ the GNSS position. The problem is
$(\hat{\mathbf{R}},\hat{\mathbf{t}})=\arg\min_{\mathbf{R}\in SO(3),\,\mathbf{t}\in\mathbb{R}^{3}}\sum_{i=1}^{M}\bigl\|\mathbf{g}_{i}-\mathbf{R}\mathbf{p}_{i}-\mathbf{t}\bigr\|_{\mathbf{W}_{i}}^{2}.$ (14)
[…]
@@ III-C3 Global Merging Using ICP on Overlapping Regions
Final map integration is performed by aligning adjacent loops using ICP on overlapping point cloud regions. Fast Point Feature Histograms (FPFH) guide initial correspondence estimation, while Octree-based indexing identifies candidate overlaps. To address ICP sensitivity to initialization and local minima, we do not start ICP from an identity guess. Instead, the initial inter-loop transform $\mathbf{T}_{0}$ is obtained from the GNSS-anchored loop frames after Sec. (GNSS-Based Loop Calibration), which provides a bounded prior for the relative pose. We further restrict ICP to verified overlap windows and use a robust point-to-plane trimmed objective. Specifically, we apply (i) a maximum correspondence distance of $1.0$ m, (ii) a normal-consistency gate with $\angle(\mathbf{n}_{a},\mathbf{n}_{b})<30^{\circ}$, and (iii) trimming ratio $\rho=0.7$ that discards the highest $30\%$ point-pair residuals. A merge is accepted only if the inlier ratio exceeds $0.30$ and the mean inlier residual is below $0.15$ m. If these criteria are not met, we re-initialize using a robust global registration solver (Go-ICP [56]) on the FPFH correspondences and then rerun the refinement. Merges that still fail are rejected to prevent contaminating the global map.
The alignment process minimizes the mean squared error between corresponding point pairs $(\mathbf{a}_{k},\mathbf{b}_{k})$ from each loop. This is formalized by the cost function $E(\mathbf{R},\mathbf{t})=\frac{1}{K}\sum_{k=1}^{K}\left\|\mathbf{R}\mathbf{a}_{k}+\mathbf{t}-\mathbf{b}_{k}\right\|^{2}$, where $\mathbf{R}$ and $\mathbf{t}$ denote the rotation and translation applied to the source point cloud. To improve robustness in dynamic environments (e.g., moving vehicles, pedestrians), a Trimmed ICP variant discards high-residual point pairs. Post-alignment, the map is refined using voxel grid downsampling (0.05 m resolution) and statistical outlier removal, enhancing compactness and structural clarity. The final map provides a reliable basis for downstream use.
@@ III-D1 Extraction of Road Centrelines
Figure 8: Velocity profiles from IMU (top, blue) and LiDAR (bottom, red) before DTW. Shaded regions indicate detected stationary phases. Misalignment across modalities underscores the need for temporal correction.
@@ III-D3 Intersection-Based Local Geometric Consistency Metrics
$\mathcal{E}_{\text{spatial}}=\frac{1}{K}\sum_{i=1}^{K}\left[\alpha\|p_{i}-m_{i}\|_{2}+\beta\cdot d_{\mathcal{T}}(p_{i},m_{i})\right]$ (16)
@@
$\mathcal{E}_{\text{env}}=\frac{1}{K}\sum_{i=1}^{K}\max_{x\in\mathcal{B}(p_{i},r)}\min_{y\in\mathcal{B}(m_{i},r)}\|x-y\|_{2}$ (17)
@@ IV Results
This section evaluates the performance of the proposed framework through a five-stage analysis. First, the experimental setup is outlined to support reproducibility. Second, sensor synchronization accuracy is examined, with emphasis on temporal alignment across LiDAR, GNSS, and IMU streams. Third, the results of 3D map reconstruction for Perth CBD are benchmarked against LiDAR-only baselines. Fourth, skeleton-based centreline metrics are employed to assess spatial accuracy and geometric consistency. Fifth, we report the end-to-end runtime and a per-stage breakdown in the Timing subsection. Each stage is described in detail in the subsequent subsections.
@@ IV-A Experimental Setup
The experiments were conducted on a high-performance workstation equipped with an Intel Xeon Gold 6330 CPU (28 cores), 512 GB of RAM, and an NVIDIA RTX 4090 GPU, running Ubuntu 20.04 LTS. To ensure reproducibility, all software components were containerized using Docker. The data processing pipeline utilized the Ouster SDK v3.1.1 for LiDAR decoding, Open3D v0.17 for point cloud filtering, and the Inertial Sense uINS SDK v2.15 for GNSS–IMU data handling. Pose graph optimization was performed using GTSAM 4.2, while NDT and ICP registration were implemented via the Point Cloud Library (PCL) v1.13.
On the 2025 Perth–CBD run (21 loops), end-to-end processing excluding acquisition took 12 h 44 m on our workstation; the synchronization step accounted for 243 m, loop registration 195 m, GNSS anchoring 126 m, and loop merging 131 m. Because our target is offline high-fidelity city-scale mapping rather than real-time deployment, we neither optimize for latency or memory nor derive formal time- or space-complexity bounds beyond the DTW order and this framing keeps the contribution centered on accurate map construction.
@@ IV-B Effectiveness of Multimodal Sensor Data Synchronization
We evaluated the proposed synchronization framework by comparing IMU and LiDAR velocity profiles before and after alignment. Accurate temporal calibration is critical for sensor fusion, especially for pose estimation and 3D mapping. Despite hardware-level synchronization via PPS triggers, residual time offsets often persist due to jitter, latency, or clock drift, which results in misaligned motion events, most evident in dynamic urban environments. As shown in Figure 8, pre-alignment profiles display mismatched transitions and inconsistent stationary phases, degrading point cloud registration and pose accuracy. To address this, we applied a DTW-based velocity matching method, which automatically estimated a $1.45\,\mathrm{s}$ offset. Post-synchronization results (Figure 9) demonstrate close alignment of both static and dynamic phases, effectively synchronizing acceleration and deceleration events across sensors.
Quantifying residuals and comparing to a constant-shift baseline. With PPS-only configuration, the median absolute timing error was 18.7 ms (p95 42.3 ms). A constant shift estimated via cross-correlation reduced the speed-profile RMSE from 0.41 to 0.22 m/s and increased the Pearson correlation from 0.72 to 0.85. DTW further reduced RMSE to 0.15 m/s and increased correlation to 0.93. The p95 residual timing drift, computed from the inferred mapping $t\mapsto t+\delta(t)$, decreased from 42.3 ms (PPS-only) to 21.4 ms (constant shift) and to 7.8 ms (DTW). Table I summarizes results over all 21 sequences. At the mapping level, replacing DTW with a constant shift increased the average centreline RMSE from 1.24 m to 1.38 m.
TABLE I: Synchronization accuracy across 21 sequences. Corr: Pearson correlation between speed profiles. RMSE: speed error. p95 drift: 95th percentile of $|\delta(t)-\mathrm{median}(\delta)|$.
Method Corr $\uparrow$ RMSE (m/s) $\downarrow$ p95 drift (ms) $\downarrow$
PPS only (no correction) 0.72 0.41 42.3
Constant shift (XCorr) 0.85 0.22 21.4
DTW (ours) 0.93 0.15 7.8
@@ IV-D1 Visual Comparison of Mapping Results
TABLE II: Comparison of road centreline alignment metrics across all test sequences for the LiDAR-only baseline and the proposed framework. Relative improvements indicate percentage reductions in alignment error.
@@ IV-D2 Loop-Based Centreline Comparison
To assess route-level consistency, we selected six representative loops from the twenty-one sequences in the Perth CBD (Seq 05, 11, 14, 16, 19, and 20). These loops span long straight segments, curved roads, and dense junctions. Figure 11 consolidates the visual evidence: for each loop, the top row shows the LiDAR-only baseline overlaid on satellite imagery, the middle row shows the reconstruction produced by LIGMA, and the bottom row presents the corresponding road centreline comparison against a 2D cartographic reference. The baseline typically exhibits drift-induced shear and irregular loop closures, with noticeable misalignment at intersections and along extended segments. In contrast, LIGMA yields coherent loop geometry and improved intersection alignment, with centrelines closely following the reference. While Fig. 11 illustrates typical behavior on six loops, the quantitative evaluation over all twenty-one sequences is summarized in Table II: our framework reduces centreline RMSE from $3.32$ m (LiDAR-only) to $1.24$ m on average, with consistent gains in the minimum, maximum, and standard deviation of the alignment error.
TABLE III: Intersection-level alignment metrics for 15 of 51 intersections (subset shown due to space constraints). CI: Centroidal Offset (m), MOD: Mean Orientation Deviation (°). Abs. = absolute reduction; Pct. = percentage reduction relative to baseline.
@@
IV-D3 Intersection-Based Alignment Evaluation
Table III summarizes results for 15 selected intersections, with percentage improvements shown in the right most columns. Table IV summarizes average centerline and intersection alignment metrics across all test sequences. The proposed approach reduces centroid offset by 61.5%, orientation deviation by 55.1%, and LEE by 59.4% over the baseline. These improvements are most pronounced at challenging junctions and map boundaries. Accurate intersection alignment supports critical tasks such as trajectory planning, and robust localization. These results demonstrate the effectiveness of the proposed GNSS-constrained fusion framework.
TABLE IV: Summary statistics of centreline and intersection alignment metrics averaged over all test sequences. CI: Centroidal Offset (m); MOD: Mean Orientation Deviation (°); RMSE: Root Mean Square Error of centreline deviation (m); SD: Standard Deviation (m). Relative improvements are computed with respect to the LiDAR-only baseline.
@@ IV-D3 Intersection-Based Alignment Evaluation
IV-E Timing
The complete 3D map construction took 14 hours and 39 minutes (879 minutes). The complete 21 loops data was collected in one go and took 115 minutes. Preprocessing took 68 minutes. The DTW synchronization comprising seeding, bounded DTW on speed/heading with anchors took 11.57 min/loop. Timestamp range trimming after synchronization took less than a minute. Loop registration using NDT and pose-graph optimization took 9.3 min/loop . GNSS-based loop calibration using weighted rigid alignment took 6 min/loop and finally ICP on overlapping regions combined with pruning and consolidation took 6.24 min/loop. The total time per loop comes out to be 41.87 minutes. We emphasize that our work targets high-fidelity city-scale 3D map construction rather than real-time deployment. Constructing an accurate 3D city map spanning 4.2 km2 in under 15 hours (including data capture) remains a significant achievement.